OpenStudy (anonymous):

What is the solution to the equation 7 −3x ≈ 9?

OpenStudy (anonymous):

7^-3x ≈ 9

OpenStudy (anonymous):

ln both sides

OpenStudy (anonymous):

Solving for x?

OpenStudy (anonymous):

\[-3x=\ln9/\ln7 \]

OpenStudy (anonymous):

yes for x

OpenStudy (anonymous):

i'm having some trouble finding this out

OpenStudy (anonymous):

\[\ln 7^(-3x) = \ln 9 \]

OpenStudy (anonymous):

i'm not sure what to do

OpenStudy (anonymous):

is it one of these? x = −0.295 x = −0.376 @ylu21

OpenStudy (anonymous):

Sorry to get back that late, how did you find the first one? the second one seems right to me

OpenStudy (anonymous):

@math871 x=(ln9/ln7)/-3=-0.376...